OpenStudy (anonymous):

How do you write an equation of the line that passes through the points: (3,5) (0,5)?

OpenStudy (anonymous):

A line has a constant slope.

OpenStudy (anonymous):

So it means that \[ \frac{y-5}{x-3} = \frac{5-5}{0-3} \]

OpenStudy (anonymous):

You have to solve for \(y\).

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

What about for (-3,-1) (0,-1)